OpenStudy (babyslapmafro):

Please help me get started on the following trigonometric substitution problem (click to see).

OpenStudy (babyslapmafro):

\[\int\limits_{}^{}e^x \sqrt{1-e ^{2x}}dx\]

OpenStudy (babyslapmafro):

How do I write the given integral as one of the following... x=asin(theta) x=atan(theta) x=asec(theta)

OpenStudy (anonymous):

Try getting rid of the exponential with a substiution u=...

OpenStudy (babyslapmafro):

ok ill try that

OpenStudy (babyslapmafro):

\[\int\limits_{}^{}u \sqrt{1-u^2}\frac{ du }{ e^x }\]

OpenStudy (babyslapmafro):

@xavier

OpenStudy (babyslapmafro):

or do i do this... \[\int\limits_{}^{}e^x \sqrt{1-u^2}\frac{ du }{ e^x }\]

OpenStudy (anonymous):

Don't turn that e^x in front of your radical into a u. Then it cancels nicely. Yes like that

OpenStudy (babyslapmafro):

ok then... \[e^x=\sin \theta\] and i plug that back into the original integral?

OpenStudy (anonymous):

Just work with the u's for now. Cancel out the e^x in front of the radical with the one under the du. Then you have a new integral which you can treat by itself forgetting all about the e^x. Then you will come back at the end and substitute back.

OpenStudy (babyslapmafro):

ok i'll show you my result when i get there

OpenStudy (anonymous):

Good luck

OpenStudy (babyslapmafro):

but when i plug sin(theta) into the equation i am integrating with respect to theta and i don't have any u's left...how do i know where to plug the e^x back in?

OpenStudy (babyslapmafro):

\[\sqrt{1-\sin ^{2}\theta}\] becomes \[\sqrt{\cos ^{2}\theta}\]

OpenStudy (babyslapmafro):

and i quickly lose track of sin(theta)

OpenStudy (anonymous):

Thats what you want to happen. Because when you do u = sin(theta) you have du = cos(theta) d(theta). Give it a try and see how it cancels

OpenStudy (babyslapmafro):

ok i integrated on this is the result \[\theta+\sin(2\theta)+C\]

OpenStudy (babyslapmafro):

and* this is the result

OpenStudy (anonymous):

Something went wrong integral of sqrt(1-u^2)du should have turned to integral of cos(theta) d(theta) Check the substiutiton one more time

OpenStudy (babyslapmafro):

du=cos(theta) d(theta)

OpenStudy (anonymous):

Ah sorry. You end up with the integral of cos^2(theta). Are you sure your answer isn't missing a factor of 1/2 along the way or a bit more. I'm assuming you used power reduction identity

OpenStudy (babyslapmafro):

w=2(theta) (1/2)w+sin(w)+c = (theta)+sin(2theta)+c

OpenStudy (babyslapmafro):

and that is after \[\int\limits_{}^{}\cos^2(\theta) d \theta=\frac{ 1 }{ 2 }\int\limits_{}^{}(1+\cos(2\theta)d \theta\]

OpenStudy (anonymous):

Let me call theta x here for sake of less typing So now you have 1/2x + 1/4sin(2x) + C.

OpenStudy (babyslapmafro):

it's not (1/2)sin(2x)?

OpenStudy (anonymous):

But you have the 1/2 outside already

OpenStudy (babyslapmafro):

where did you get (1/4)

OpenStudy (anonymous):

What is the integral of just cos(2x)?

OpenStudy (babyslapmafro):

oh ok i see now